OpenStudy (anonymous):

Ramon bought two pullovers on a buy one, get one 30% off offer. Each pullover had a list price of $20. how much did Ramon pay for the pullovers?

OpenStudy (ja1):

Okay so to find the discount we multiply the rate by the price. So: \[30 * 20 = ?\]

OpenStudy (ja1):

Wait no, \[.30 * 20 = \]

OpenStudy (ja1):

So just plug that into a calculator :)

OpenStudy (anonymous):

are you sure because im getting 6 but thats not one of my answer choices

OpenStudy (ja1):

nono that is just the first part ::)

OpenStudy (ja1):

Now you subtract the discount by the original price.

OpenStudy (ja1):

So: \[20 - 6\]

OpenStudy (anonymous):

thats 14 and what do i do next

OpenStudy (ja1):

So if the original pullover is 20 dollars and the other one is 14 then: \[20 + 14 = ?\]

OpenStudy (anonymous):

34

OpenStudy (ja1):

There you go :)

OpenStudy (anonymous):

thanks you were alot of help :)

OpenStudy (ja1):

No problem :)